Help I hate setting up word problems I just cant master it the question is:
find four consecutive integers such that the product of the two largest is 46 more than the product of the two smallest integers.
Thanks

Answer by checkley77(12844)

Let x, x+1, x+2 & x+3 be the four integers.
(x+2)(x+3)=x(x+1)+46
x^2+5x+6=x^2+x+46
x^2-x^2+5x-x=46-6
4x=40
x=40/4
x=10 for the smallest integer.
10=1=11 for the next integer.
10=2=12 for the third integer.
10=3=13 for the largest integer.
12*13=10*11+46
156=110+46
156=156
Hope this helps.